A tourist, who weighs 805 N, is walking through the woods and crosses a small horizontal bridge. The bridge rests on two concrete supports, one at each end. He stops one-fourth of the way along the bridge. Assume that the board of the bridge has negligible weight. What is the magnitude of the vertical force that a concrete support exerts on the bridge at the far end?


1
Expert's answer
2021-04-13T06:49:34-0400

"R_1" and "R_2" are the vertical force at near and far end respectively.


Taking moment at point A;

"-R_1×0-w×\frac {L} {4}+R_2×L=0"


"=R_2L=W\frac {L} {4}"


"R_2=\frac {w} {4}=\frac {805}{4}"


"R_2=201.25N"


Thus the vertical force acting on the far end is 201.25N